OpenStudy (anonymous):

how do you factor 3x^2+3-60??

OpenStudy (anonymous):

to factor 3x2 − 3x − 60 you have to divide the whole equation by 3 because x^2 cannot be 3x^2 x^2-x-20 cannot be factored so you would have to use quadratic formula -b+-√b^2-4ac ---------------- 2 1+-√1^2-4(1)(-20)=1+-√1+80=1+-√81=9 1+9/2=5 or 1-9/2=-4 (x+5)(x-4)

OpenStudy (anonymous):

Thanks sooo much!

OpenStudy (anonymous):

ms rry 4 dis/// but y z no buddy givin meddal!!